The Royal Fruit Company produces two types of fruit drinks. The first type is 60% pure fruit juice, and the second type is 85% pure fruit juice. The company is attempting to produce a fruit drink that contains 75% pure fruit juice. How many pints of each of the two existing types of drink must be used to make 30 pints of a mixture that is 75% pure fruit juice?
Answer by ankor@dixie-net.com(22740)

The Royal Fruit Company produces two types of fruit drinks.
The first type is 60% pure fruit juice,
and the second type is 85% pure fruit juice.
The company is attempting to produce a fruit drink that contains 75% pure fruit juice.
How many pints of each of the two existing types of drink must be used to make 30 pints of a mixture that is 75% pure fruit juice?
:
let x = amt of 85% juice
the resulting amt is to be 30 pints, therefore
(30-x) = amt of 60% juice
:
The decimal equiv equation
.85x + .60(30 - x) = .75(30)
.85x + 18 - .60x = 22.5
.85x - .60x = 22.5 - 18
.25x = 4.5
x = 4.5/.25
x = 18 pints of 85% juice
and
30 - 18 = 12 Pints of 60% juice
:
:
Check this
.85(18) + .60(12) = .75(30)
15.3 + 7.2 = 22.5